OpenStudy (daenio):

8. Jimmy really wants to move out of his cheap apartment. He buys some lottery tickets from the local 7-11. The probability of winning on a ticket is 1 in 10. a) What is the probability Jimmy will win on his second ticket?

OpenStudy (daenio):

@blockcolder

OpenStudy (daenio):

@Mertsj @amistre64 @joemath314159

OpenStudy (amistre64):

talk about yelling for attention

OpenStudy (daenio):

LOL, I just really want help with some questions.

OpenStudy (amistre64):

the probability seems to be per ticket; so maybe 1/10 ? i cant be sure becasue i tend to misread alot of these prob things

OpenStudy (daenio):

It's a binomial distribution question so I guess a formula is required.

OpenStudy (amistre64):

whats the probability of getting a tails on your second coin toss?

OpenStudy (amistre64):

since each toss, or each ticket,has the same chance of winning; id say the prob stays put.

OpenStudy (daenio):

Oh! I think it would be like this: 1 in 10 translates into 10%, which is 0.10. q=1-p =1-0.10 =0.90 P(x)=q^x p =0.90^2 (0.10) =0.81 (0.10) =0.081

OpenStudy (amistre64):

it might, but i just cant read that much detail from the question to be sure